# Arboretum: A Large Multimodal Dataset Enabling AI for Biodiversity

### v1

@@ Appendix B ArborCLIP training details
Appendix C Additional ArborCLIP results
In Table 7, we report model performance at different levels of the taxonomic hierarchy. Generally, we find that models trained on web-scraped data perform better with common names, whereas models trained on specialist datasets perform better when using scientific names. Additionally, models trained on web-scraped data excel at classifying at the highest taxonomic level (kingdom), while models begin to benefit from specialist datasets like Arboretum-40M and Tree-of-Life-10M at the lower taxonomic levels (order and species).
However, ArborCLIP shows a performance decline at taxonomic levels below the species level. This is likely because our training metadata structure allows for classifications solely by referring to species information. From a practical standpoint, this is not problematic for the species in our test set since ArborCLIP is highly accurate at the species level, and higher-level taxa can be deterministically derived from the lower ones.
Furthermore, the OpenCLIP and MetaCLIP baselines outperform ArborCLIP on the life stages benchmark. This highlights the importance of retaining the general linguistic capabilities of the pretrained CLIP models for hybrid tasks.
Table 7: Different levels of taxonomic classification reveal the strengths and weaknesses of different models. DS stands for dataset, Arch for model architecture AU stands for Arboretum-Unseen, AB stands for Arboretum-Balanced, BCR for BioCLIP-Rare, K for Kingdom taxonomy, O for Order taxonomy, and CN for species taxonomy using common names.
DS Arch AU-K AU-O AU-CN AB-K AB-O AB-CN BCR-CN
OpenAI ViT-B-16 78 26.7 18.6 79.9 26.7 13.1 31.6
BioCLIP ViT-B-16 75.5 38.4 41.8 73.2 34.5 42.7 32.1
MetaCLIP-cc ViT-L-14 90.3 33.2 27.9 87.1 34.5 19.9 42.4
ArborCLIP from OpenAI ViT-B-16 62.9 13.5 29.7 56.5 12.4 70.8 22.2
ArborCLIP from BioCLIP ViT-B-16 62.8 17.8 24 65.7 14.5 52.6 23.6
ArborCLIP from MetaCLIP ViT-L-14 68.3 9.6 4.3 58.1 3.8 15.7 8.8

### v2

@@ A.8 BioTrove-CLIP training details
A.9 Additional BioTrove-CLIP results
In Table 8, we report model performance at different levels of the taxonomic hierarchy. Generally, we find that models trained on web-scraped data perform better with common names, whereas models trained on specialist datasets perform better when using scientific names. Additionally, models trained on web-scraped data excel at classifying at the highest taxonomic level (kingdom), while models begin to benefit from specialist datasets like BioTrove-Train and Tree-of-Life-10M at the lower taxonomic levels (order and species).
However, BioTrove-CLIP shows a performance decline at taxonomic levels below the species level. This is likely because our training metadata structure allows for classifications solely by referring to species information. From a practical standpoint, this is not problematic for the species in our test set since BioTrove-CLIP is highly accurate at the species level, and higher-level taxa can be deterministically derived from the lower ones.
Furthermore, the OpenCLIP and MetaCLIP baselines outperform BioTrove-CLIP on the life stages benchmark. This highlights the importance of retaining the general linguistic capabilities of the pretrained CLIP models for hybrid tasks.
Table 8: Performance Comparison Across Benchmarks: This table compares the performance of BC-iNat21 (trained solely on the iNaturalist 2021 dataset) and BT-Clip (trained from the BioCLIP checkpoint, originally trained on the TreeOfLife dataset). Metrics include Top-1 Accuracy and Top-5 Accuracy.
Benchmark BC-iNat21 Top-1 BC-iNat21 Top-5 BT-Clip Top-1 BT-Clip Top-5
BioTrove Unseen 0.2100 0.3470 0.5380 0.8220
Fungi 0.4420 0.7550 0.5390 0.7590
Life-Stages 0.2867 0.8617 0.1500 0.8600
DeepWeeds 0.2057 0.6897 0.1840 0.5740
Insects-2 0.0103 0.0483 0.1690 0.5710
Birds-525 0.5030 0.6330 0.5710 0.7540
BioCLIP-Rare 0.1490 0.2790 0.2370 0.7600
BioTrove Balanced 0.5020 0.6450 0.5180 0.6610
A.10 Additional BioTrove-CLIP Comparative Analysis
We conducted a comparative evaluation of the top-1 and top-5 zero-shot accuracy of the BioCLIP model, which was trained exclusively on the iNaturalist 2021 (iNat21) dataset, and the BioTrove-CLIP model, initialized from BioCLIP checkpoints originally trained on the TreeOfLife dataset. The comparison highlights the performance differences across various benchmarks, as presented in Table 9.
Our analysis shows that models trained on the BioTrove dataset consistently outperform those trained solely on iNat21, particularly in benchmarks such as BioTrove-Unseen, Fungi, and Insects-2. While certain benchmarks like Life-Stages and DeepWeeds show moderate differences, the results emphasize the advantages of training on BioTrove, leading to enhanced model accuracy and robustness.
The following table provides detailed performance metrics for both models across various benchmarks, comparing their top-1 and top-5 accuracy scores with associated confidence intervals.
Table 9: Performance Comparison Across Benchmarks: This table compares the performance of BC-iNat21 (trained solely on the iNaturalist 2021 dataset) and BT-Clip (trained from the BioCLIP checkpoint, originally trained on the TreeOfLife dataset). Metrics include Top-1 Accuracy and Top-5 Accuracy. BC-iNat21 refers to BioCLIP (iNat21), and BT-Clip refers to BioTrove-CLIP (BioCLIP checkpoint from TreeOfLife).
Benchmark BC-iNat21 Top-1 Acc. BC-iNat21 Top-5 Acc. BT-Clip Top-1 Acc. BT-Clip Top-5 Acc.
BioTrove-Unseen 0.2100 0.3470 0.5380 0.8220
Fungi 0.4420 0.7550 0.5390 0.7590
Life-Stages 0.2867 0.8617 0.1500 0.8600
DeepWeeds 0.2057 0.6897 0.1840 0.5740
Insects-2 0.0103 0.0483 0.1690 0.5710
Birds-525 0.5030 0.6330 0.5710 0.7540
BioCLIP-Rare 0.1490 0.2790 0.2370 0.7600
BioTrove Balanced 0.5020 0.6450 0.5180 0.6610
As demonstrated, the model trained on BioTrove exhibits superior performance in most categories, particularly when evaluated on rare and unseen species, underscoring the importance of diverse and large-scale datasets like BioTrove for enhancing biodiversity AI models.